OpenStudy (anonymous):

if f(x)=4/x-4 and g(x)= 14/x+8 what is f/g

OpenStudy (anonymous):

invert and multiply

OpenStudy (anonymous):

\[\frac{\frac{4}{x-4}}{\frac{14}{x+8}}=\frac{4}{x-4}\times \frac{x+8}{14}\]

OpenStudy (anonymous):

ans didnt work

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

how do you type all real numbers except x=8 mathematically